[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Nov 03, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Nov 3 11:35:51 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The grunt and cumulus N-AWIPS workstations did not come back up after a  
reboot late yesterday. Dave called the SDM and they assigned a sys admin  
from College Park to bring them back up, which was completed within about  
an hour after they went down. It was the same issue with the video drivers  
not loading properly. Mark will check with NCO today to determine why the  
procedure to automatically load the drivers failed.


HSU:
One of the ATCF workstations was slow yesterday. That workstation had too  
many instances of the ATCF opened at once. Closing some of those resolved  
the problem.
Other items:
The LDM on the logger server has remained stable. Dave is working on a  
program to automatically restart that LDM if it gets hung up. That should  
be implemented next week.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161103/8d394bf5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list